OpenStudy (anonymous):

Complete the following ordered pairs for the equation. x + 4y = 10 (-2, _) (_, 2) (6, _)

OpenStudy (aroub):

ok,first one (-2,2) second one (2,2) last one (6,1)

OpenStudy (anonymous):

thanks aroub

OpenStudy (aroub):

anytime mathquiz! :D

OpenStudy (anonymous):

Plug the number into the equation respectively. #1 -2+4y=10 y=3 so it's (-2,3), #2 x+4*2=10 x=2 so it's(2,2) #3 6+4y=10 y=1 so it's(6,1), I guess the first answer of aroub is wrong...